How do i solve for X in the following question 15x-2=21x+10
Found 3 solutions by haileytucki, stanbon, nyc_function:
Answer by haileytucki(390)

15x-2=21x+10
Since 21x contains the variable to solve for, move it to the left-hand side of the equation by subtracting 21x from both sides.
15x-2-21x=10
Since 15x and -21x are like terms, add -21x to 15x to get -6x.
-6x-2=10
Since -2 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 2 to both sides.
-6x=2+10
Add 10 to 2 to get 12.
-6x=12
Divide each term in the equation by -6.
-(6x)/(-6)=(12)/(-6)
Simplify the left-hand side of the equation by canceling the common factors.
x=(12)/(-6)
Simplify the right-hand side of the equation by simplifying each term.
x=-2

Answer by stanbon(75887)

15x-2=21x+10
-----
Subtract 15x from both sides.
Subtract 10 from both sides.
-12 = 6x
x = -2
-------------
Cheers,
Stan H.

Answer by nyc_function(2741)

15x - 2 = 21x + 10

All you have to do is combine like terms.


Don't forget to change the signs when crossing the equal sign.


15x - 21x = 2 + 10

-6x = 12

x = 12/-6

x = -2

THE PROOF:

Let x = -2

15(-2)-2 = 21(-2) + 10

-30 - 2 = -42 + 10

-32 = -32...It checks!!

We now know that x = -2 is right.